# Logical perspectives on learning statistical objects

### v1

@@ Appendix A Agnostic online learnabilty and duality: Proof of Proposition 1
Clearly, if the roles of parameters are swapped, the same thing holds.
Appendix B Continuous logic over classical first-order structures: proof of Proposition 2
Recall that we allow continuous logic to be defined over a classical first-order structure $\mathfrak{M}$. We now have more formulas, including infinite convergent sums and applying continuous functions. For example, if our structure has infinitely many unary predicates $U_{i}(x)$, we can form a new formula
$\chi(x)=\Sigma_{i}\frac{U_{i}(x)}{2^{i}}.$
If we have only a single binary predicate $G(x,y)$, we could similarly start by letting $\phi_{i}(x)$ state that there is a $G$ path of length $i$ originating at $x$, and then set
$\chi(x)=\Sigma_{i}\frac{\phi_{i}(x)}{2^{i}}.$
Recall Proposition 2 from the body of the paper:
If a first-order structure $\mathfrak{M}$ is NIP, then for any general continuous logic partitioned formula $\phi$, ${\mathcal{H}}_{\phi}$ has finite fat-shattering dimension and is thus agnostic PAC learnable. Thus $\mathfrak{C}(\mathfrak{M})$ is also NIP.
That is, if $\mathfrak{M}$ is NIP (equivalently, all partitioned formulas are agnostic PAC learnable) as a classical first-order structure, then it is still NIP as a CL structure: even though we have more formulas, hence more hypothesis classes, the new ones are still learnable.
This is implicit in Ben Yaacov (2009), but we spell out a few more details here. Recall the distinction between basic formulas and general formulas of CL: the latter are formed by closing under convergent sum. We first note that general formulas can be uniformly approximated by basic formulas. From this it follows that to show general partitioned formulas are NIP (i.e. induce function classes with finite $\gamma$ fat-shattering dimension), it suffices to show the same for basic formulas.
Call a CL formula essentially FO if it is of the form $\Sigma_{i\leq n}\chi_{i}(\vec{x})\cdot r_{i}$ where $\chi_{i}$ is the characteristic function of a first-order formula. It is easy to see (and see (Ben Yaacov et al., 2008, Remark 9.21) for a proof) that basic formulas are essentially FO.
If a classical structure is NIP then each $\chi_{i}$ has finite VC dimension, hence trivially each one has finite $\gamma$ fat-shattering dimension. It is easy to see that finite fat-shattering dimension is preserved under scalar multiples and summing.
Appendix C Lemmas on continuous logic, NIP, and definable functions: proof of Lemma 1
Our goal will be to prove Lemma 1 from the body. Recall that we deal there with a first-order structure $\mathfrak{M}$ over the reals that has at least the ordering relation. We first note the following:
Lemma 13.
The function $f_{[0,1]}:{\mathcal{R}}\to[0,1]$ given by
$f(x)=\begin{cases}0&x\leq 0\\ x&0\leq x\leq 1\\ 1&1\leq x\end{cases}$
is a formula in $\mathfrak{C}(\mathfrak{M})$.
Proof.
For each $n$, let $\phi_{n}$ be the indicator function of the set $\bigcup_{i=1}^{2^{n-1}}\left(\frac{2i-1}{2^{n}},\frac{2i}{2^{n}}\right]$. Each of these is the indicator function of a definable set in $\mathfrak{M}$, so it is a formula in $\mathfrak{C}(\mathfrak{M})$. Also let $\psi$ be the indicator function for the definable set $(1,\infty)$. Then we can take an infinite weighted sum, and see that $\psi+\sum_{n=1}^{\infty}\frac{\phi_{n}}{2^{n}}$ is also a formula. At any $x\in[0,1]$, this sum will evaluate to $x$, as $\phi_{n}(x)$ will give the $n^{th}$ bit in a binary expansion of $x$. This will also evaluate to 00 for $x<0$, and to $\psi(x)=1$ for $x>1$, so this sum formula is $f_{[0,1]}$. ∎
Once we have this formula, we are able to construct all other bounded definable functions.
Recall the statement of Lemma 1:
Let $f:{\mathcal{R}}^{n}\to[0,1]$ be a definable function in $\mathfrak{M}$. Then it is also a formula in continuous logic over $\mathfrak{C}(\mathfrak{M})$.
Proof.
We may apply any formula in one variable to a definable function and get another formula, and in this case, we find that $f_{[0,1]}\circ f=f$, so $f$ is a formula. ∎
Appendix D Proof of Proposition 3: inducing measures
Recall the statement of the proposition from the body of the paper:
For any range $X$, we can choose $(\Omega,\Sigma,\mu)$ and well-behaved $RV_{X}$ such that for each measure $\mu^{\prime},\Sigma^{\prime}$ on $X$ there is an $F\in RV_{X}$ that induces $\mu^{\prime}$ from $\mu$.
Proof.
We choose ${\mathcal{P}}$ to be a product of all measure spaces on $X$. Then every measure is induced as a projection, and we take $RV_{X}$ to be all such projections.

### v2

@@ Appendix A Agnostic online learnabilty and duality: Proof of Proposition 3
Clearly, if the roles of range and parameter space are swapped, the same thing holds.
Appendix B Justifying and generalizing Example 4
Recall that in the body of the paper we presented Example 4, an example of a hypothesis class of functions. We claimed that this was agnostic PAC learnable. This can be argued directly via computation of fat-shattering dimensions. Here we give a more general argument, deriving from logic.
Consider the real numbers with operations addition, multiplication, and the inequality: this is a real-closed order field. We recall what it means for a set of tuples in the reals to be first-order definable over this vocabulary The terms of first-order logic are built up from real constants and variables by applying the functios addition and multiplication: that is, they are polynomials with real coefficients. The atomic formulas are inequalities between terms. The formulas of first-order logic are built up from atomic formulas via the Boolean operations ,,$\wedge$,$\vee$,$\neg$ , ∨ , ¬ and the quantifiers $\exists x$, $\forall x$.
A partitioned formula $\phi(x_{1}\ldots x_{j};w_{1}\ldots w_{\ell})$ is a formula where the free variables are partitioned into two subsets, $\vec{x}$ and $\vec{y}$. Such a formula is associated with a family of subsets of ${\mathcal{R}}^{j}$, indexed by ${\mathcal{R}}^{\ell}$, in the obvious way. We can similarly talk about a formula with a partition of the free variables into three parts.
A $3$-partitioned formula of the form $\phi(x_{1}\ldots x_{j};y_{1}\ldots y_{k};w_{1}\ldots w_{\ell})$ is a definable family of functions if for every $\vec{w}$ and $\vec{x}$, there is exactly one $\vec{y}$ that makes the formula true. Such a formula is associated with a family of functions from ${\mathcal{R}}^{j}$ to ${\mathcal{R}}^{k}$, indexed by ${\mathcal{R}}^{\ell}$.
Note that the class of Example 4 is of this form.
The following proposition follows from the fact that the real ordered field is an “NIP structure”:
Proposition 62
Every definable family of functions over the real ordered field has finite $\gamma$ fat-shattering dimension for each $\gamma>0$, and is thus agnostic PAC learnable.
Although this is also probably well known, we sketch a proof:
Proof  To use the terminology within model theory, the real-ordered field is o-minimal, and for every o-minimal structure over the reals, it is shown in van den Dries (1998) that every definable family family of subsets in an o-minimal structure has finite VC dimension. Thus if we have $\phi(\vec{x};\vec{w})$, with $\phi$ a formula over the real field, the hypothesis class consisting of functions $h_{\vec{w}}$, the characteristic function $\phi(\vec{x};\vec{w})$ has finite $\gamma$ fat-shattering dimension for each $\gamma$. Thus the same holds for finite sums (with fixed number of summands) of such characteristic functions. A definable family of functions, like the family in Example 4, can be approximated within any tolerance in the sup norm, uniformly in the parameters $\vec{w}$ by finite sums of characteristic functions. Thus the proposition holds.
Appendix C Measurability issues
In the body of this paper, there are two important steps where our arguments ignored some issues related to measurability assumptions, and two points at which we cite other work without clarifying the measurability assumptions. We spell out these steps and the assumptions here.
Let us first discuss the agnostic PAC case. We have given two arguments for preservation of agnostic PAC learnability to statistical classes in the body of the paper. Both of them proceed by proving bounds on certain combinatorial dimensions, like $\gamma$ fat-shattering. Our story was that this implies finite GC-dimension, which in turn implies agnostic PAC learnability, a standard chain of deductions. However, the argument from combinatorial bounds to GC-dimension bounds in prior work requires some measurability assumptions, which are often not spelled out in detail in the literature: see, e.g. Lothar Sebastian Krapp and Laura Wirth (2025) for an overview of these issues.
We will not provide exact measurability conditions. But we will show that countability of the parameter space is sufficient for our sample complexity arguments about the distribution class, and similarly countability of the range space for the arguments about the sample complexity of the dual distribution class.
Recall that a key point of the argument is to bound the GC dimensions of the expectation class of a measurable family. We will argue that:
  • for this argument it suffices that the parameter space is separable: informally, this means we can approximate the behaviour by looking at a countable subset of the parameter space.
  • when we revisit the embedding of the distribution class as a measurable family from Section 3.2, it produces a separable measurable family as long as the parameter space of the base class is countable.
Similarly, in agnostic online case, our proof of Theorem 34 requires additional assumptions, including those required by results cited from Rakhlin et al. (2015b). We will see that these arguments also work in the context of a separable parameter space.
C.1 The Necessary Measurability Assumptions
Recall Lemma 20:
Let $(\Omega,\Sigma,\mu)$ be a probability space, and let ${\mathcal{F}}=({\mathcal{H}}_{\omega}:\omega\in\Omega)$ be a measurable family of hypothesis classes on $X$.
$0pt({\mathbb{E}}{\mathcal{F}}(\bar{x},Y),\beta)\leq\mathbb{E}_{\mu}[0pt({\mathcal{H}}_{\omega}(\bar{x},Y),\beta)].$
The first issue is that for $\mathbb{E}_{\mu}[0pt({\mathcal{H}}_{\omega}(\bar{x},Y),\vec{b})]$ in the statement of Lemma 20 to be well-defined, we need the measurable family ${\mathcal{F}}$ to satisfy a stronger measurability assumption:
Definition 63
Assume that $(\Omega,\Sigma,\mu)$ is a probability space. Say that a measurable family ${\mathcal{F}}=({\mathcal{H}}_{\omega}:\omega\in\Omega)$ of hypothesis classes on $X$ parameterized by $Y$ is strongly measurable when for each $\bar{x}\in X^{n}$ and $\vec{b}\in{\mathcal{R}}^{n}$, the function $\omega\mapsto 0pt({\mathcal{H}}_{\omega}(\bar{x},Y),\vec{b})$ is measurable.
Another place measurability is required is defining Glivenko-Cantelli dimensions. The Glivenko-Cantelli dimensions of a hypothesis class ${\mathcal{H}}$ are defined in terms of the measures of the sets
[…]
for $\varepsilon>0,m$, and a distribution $D$, which we could rewrite as the union
[…]
which in general need not be measurable, although they are if for each $m$, the following function is measurable:
[…]
In summary, when we refer to the Glivenko-Cantelli dimensions of ${\mathcal{H}}$, we implicitly assume that ${\mathcal{H}}$ has the following property:
Definition 64 (well-defined GC dimensions)
Say a hypothesis class ${\mathcal{H}}$ on $X$ has well-defined Glivenko-Cantelli dimensions if for all $\varepsilon>0$, all $m$, and all distributions $D$ on measure space on $X$ induced by ${\mathcal{H}}$, the set
[…]
is measurable with respect to the measure space on $X$ induced by ${\mathcal{H}}$.
If ${\mathcal{H}}$ has well-defined Glivenko-Cantelli dimensions, then the bounds in Fact 10 on the learnability of ${\mathcal{H}}$ hold.
The two results we cited without measurability assumptions are Facts 11 and 19:
Fact 11 is used in the proof of Theorem 15 to deduce bounds on Glivenko-Cantelli dimensions from Rademacher mean width. The result this rephrases, (Wainwright, 2019, Theorem 4.10), assumes measurability in the form of well-defined Glivenko-Cantelli dimensions. It is observed at the beginning of (Wainwright, 2019, Section 4.4) that this measurability assumption holds in the case of a parameter space which is either countable or, in an appropriate metric, separable.
Fact 19 is used in the proof of Theorem 34 to deduce online learnability from sequential Rademacher mean width. It is cited from Rakhlin et al. (2015b), where $X$ and $Y$ are assumed to be separable metric spaces, with ${\mathcal{H}}:X\times Y\to[0,1]$ continuous (or at least, lower-semicontinuous). In the case where $X$ and $Y$ are countable, this can be satisfied trivially by giving each the discrete metric ($d(a,b)=1$ when $a\neq b$). We will show that this assumption is also satisfied when $X$ is countable and $Y$ is a space of random variables on a countable set.
It is worth noting that Theorem 29 and Corollary 45 provide alternate proofs that agnostic PAC and online learning respectively are preserved under passing to the distribution class. The proofs of both of these theorems simply show that finite (sequential) fat-shattering dimension is preserved, from which one can deduce that learnability is as well. However, this latter deduction is once again subject to measurability constraints. In the PAC case, well-defined Glivenko-Cantelli dimensionssuffice, and in the online case, separability suffices.
C.2 Countable Parameter Spaces
We now check that if $Y$ is countable, then these measurability assumptions are satisfied.
Lemma 65
If ${\mathcal{F}}$ is a measurable family of hypothesis classes on $X$ parameterized by $Y$, and $Y$ is countable, then ${\mathcal{F}}$ is strongly measurable.
Proof  Assume $Y$ is countable, and fix $\bar{x}=(x_{1},\dots,x_{n})\in X^{n}$, $\vec{b}=(b_{1},\dots,b_{n})\in{\mathcal{R}}^{n}$. For any $\omega\in\Omega$, we can view ${\mathcal{H}}_{\omega}$ as a function ${\mathcal{H}}_{\omega}:X\times Y\to[0,1]$. Then expanding definitionally,
$0pt({\mathcal{H}}_{\omega}(\bar{x},Y),\vec{b})=\sup_{y\in Y}\sum_{i=1}^{n}{\mathcal{H}}_{\omega}(x_{i},y)b_{i}.$
The assumption that ${\mathcal{F}}$ is measurable means that for each $x\in X,y\in Y$, the function $\omega\mapsto{\mathcal{H}}_{\omega}(x,y)$ is measurable, from which we see that for each $y\in Y$, the function $\omega\mapsto\sum_{i=1}^{n}{\mathcal{H}}_{\omega}(x_{i},y)b_{i}$ is measurable. A countable supremum of measurable functions is measurable, so $\omega\mapsto 0pt({\mathcal{H}}_{\omega}(\bar{x},Y),\vec{b})$ is measurable, making ${\mathcal{F}}$ a strongly measurable class.
It is also clear that if ${\mathcal{H}}$ is a class parameterized by a countable set $Y$, the Glivenko-Cantelli dimensions are well-defined, because the sets in question are countable unions.
We can also conclude that if ${\mathcal{F}}$ is a measurable family of hypothesis classes parameterized by countable $Y$, then Theorem 15 holds. Because a countable set is also separable in any metric, the measurability requirements for Fact 19 are satisfied, so Theorem 34 holds.
C.3 From countable parameter spaces to the parameters of random variable classes
When we analyze a class ${\mathcal{H}}$ under the assumption that its parameter set $Y$ is countable, we eventually have to consider the class ${\mathcal{H}}(\mathrm{CompatParamRV}({\mathcal{H}}))$, parameterized by $\mathrm{CompatParamRV}({\mathcal{H}})$, which need not be countable. However, we will show that it is separable in a suitable metric, and show that this suffices for the rest of our measurability requirements.
Lemma 66
If $Y$ is countable, then the set $\mathrm{CompatParamRV}({\mathcal{H}})$ is separable with the metric where
$d(Y_{1},Y_{2})=\sum_{y\in Y}\left|\mathbb{P}[Y_{1}=y]-\mathbb{P}[Y_{2}=y]\right|.$
Proof  Let $A$ be the set of all $Y^{*}\in\mathrm{CompatParamRV}({\mathcal{H}})$ such that for all $y\in Y,$ $\mathbb{P}[Y^{*}=y]\in\mathbb{Q}$, for all but finitely many $y\in Y$, $\mathbb{P}[Y^{*}=y]=0$. This set is countable, and to show that it is dense, we will show that for all $Y^{\prime}\in\mathrm{CompatParamRV}({\mathcal{H}})$, and $\varepsilon>0$, there is some $Y_{\varepsilon}\in A$ such that $d(Y^{\prime},Y_{\varepsilon})\leq\varepsilon$.
As $\sum_{y\in Y}\mathbb{P}[Y^{\prime}=y]=1$, there is some finite set $S\subseteq Y$ with $\sum_{y\in S}\mathbb{P}[Y^{\prime}=y]\geq 1-\frac{\varepsilon}{3}$. Then by density of the rational numbers, we can find nonnegative rational numbers $(r_{y}:y\in S)$ such that $\sum_{y\in S}r_{y}=1$, and $\sum_{y\in S}|\mathbb{P}[Y^{\prime}=y]-r_{y}|\leq\frac{2\varepsilon}{3}$. We then define $Y_{\varepsilon}$ so that if $y\in S$, then $\mathbb{P}[Y_{\varepsilon}=y]=r_{y}$, and otherwise, $\mathbb{P}[Y_{\varepsilon}=y]=0$.
Then $d(Y^{\prime},Y_{\varepsilon})=\sum_{y\in S}\sum_{y\in S}|\mathbb{P}[Y^{\prime}=y]-r_{y}|+\sum_{y\in Y\setminus S}|\mathbb{P}[Y^{\prime}=y]-0|\leq\frac{2\varepsilon}{3}+\frac{\varepsilon}{3}=\varepsilon.$
We now check that separability in this particular metric suffices to imply well-defined Glivenko-Cantelli dimension, justifying the application of Fact 11 in the proof of Theorem 15.
Theorem 67
If $Y$ is countable, then for any $\delta,\varepsilon>0$, $GC_{\varepsilon,\delta}\left({\mathcal{H}}(\mathrm{CompatParamRV}({\mathcal{H}}))\right)$ is well-defined.
Proof  It suffices to show that for all $m$, the function
[…]
is measurable. To do this, let $A$ be a countable dense set of $\mathrm{CompatParamRV}({\mathcal{H}})$ in the metric from Lemma 66, and we will show that the functions $f_{Y^{\prime}}$ defined by
[…]
are themselves measurable, and that $\{f_{Y^{\prime}}:Y^{\prime}\in A\}$ is dense in $\{f_{Y^{\prime}}:Y^{\prime}\in\mathrm{CompatParamRV}({\mathcal{H}})\}$, in the $\sup$-metric, so any $f_{Y^{\prime}}$ is a uniform limit of elements of $\{f_{Y^{\prime}}:Y^{\prime}\in A\}$, from which we conclude that
$\sup_{Y^{\prime}\in\mathrm{CompatParamRV}({\mathcal{H}})}f_{Y^{\prime}}=\sup_{Y^{\prime}\in A}f_{Y^{\prime}},$
and the latter countable supremum is measurable.
First, let $Y^{\prime}\in\mathrm{CompatParamRV}({\mathcal{H}})$. The function $f_{Y^{\prime}}$ is measurable, because it is the absolute value of a sum of measurable functions, as $x\mapsto\mathbb{E}_{\mu}[h_{Y^{\prime}}(x)]$ is the uniform limit of linear combinations of measurable functions $h_{y}(x)$ for $y\in Y^{\prime}$.
Then we show that $\{f_{Y^{\prime}}:Y^{\prime}\in A\}$ is dense in $\{f_{Y^{\prime}}:Y^{\prime}\in\mathrm{CompatParamRV}({\mathcal{H}})\}$, in the $\sup$-metric, by showing that if $Y_{1},Y_{2}\in\mathrm{CompatParamRV}({\mathcal{H}})$, then $\sup_{\bar{x}}|f_{Y_{1}}(\bar{x})-f_{Y_{2}}(\bar{x})|\leq 2d(Y_{1},Y_{2})$.
We see that for every $x\in X$,
$\mathbb{E}_{\mu}[\left|h_{Y_{1}}(x)-h_{Y_{2}}(x)\right|]$ $\leq\sum_{y\in Y}h_{y}(x)\left|\mathbb{P}[Y_{1}=y]-\mathbb{P}[Y_{1}=y]\right|$
$\leq\sum_{y\in Y}\left|\mathbb{P}[Y_{1}=y]-\mathbb{P}[Y_{1}=y]\right|$
$=d(Y_{1},Y_{2}),$
so for every $\bar{x}\in X^{m}$,
$f_{Y_{1}}(\bar{x})-f_{Y_{2}}(\bar{x})$
[…]
Furthermore, to show that the assumptions of Fact 19 are satisfied, we check the following:
Lemma 68
Suppose ${\mathcal{H}}$ is a hypothesis class with range space $X$ and parameter space $Y$, which are both countable. Then ${\mathcal{H}}(\mathrm{CompatParamRV}({\mathcal{H}})):X\times\mathrm{CompatParamRV}({\mathcal{H}})\to[0,1]$ is continuous, where $X$ has the discrete metric and $\mathrm{CompatParamRV}({\mathcal{H}})$ has the above separable metric.
Proof  Because $X$ has the discrete metric, it suffices to show that for each $x\in X$, the function
$f_{x}:Y^{\prime}\mapsto({\mathcal{H}}(\mathrm{CompatParamRV}({\mathcal{H}})))_{Y^{\prime}}(x)$
is continuous.
In fact, we will see that it is 1-Lipschitz. Suppose $Y_{1},Y_{2}\in\mathrm{CompatParamRV}({\mathcal{H}}))$. Then
$|f_{x}(Y_{1})-f_{x}(Y_{2})|=$ $\left|\sum_{y\in Y}\mathbb{P}[Y_{1}=y]h_{y}(x)-\sum_{y\in Y}\mathbb{P}[Y_{1}=y]h_{y}(x)\right|$
$=$ $\left|\sum_{y\in Y}\left(\mathbb{P}[Y_{1}=y]-\mathbb{P}[Y_{1}=y]\right)h_{y}(x)\right|$
$\leq$ $\sum_{y\in Y}\left|\mathbb{P}[Y_{1}=y]-\mathbb{P}[Y_{1}=y]\right|h_{y}(x)$
$\leq$ $\sum_{y\in Y}\left|\mathbb{P}[Y_{1}=y]-\mathbb{P}[Y_{1}=y]\right|$
$=d(Y_{1},Y_{2}).$
C.4 Without Countability Assumptions
If we place no countability or separability assumptions on $Y$, then we can still prove results which transfer from a base class to a statistical class, but which only deal with combinatorial dimensions, rather than learnability. Here would be one transfer result for the agnostic PAC case:
Theorem 69
If ${\mathcal{F}}$ is a measurable family of hypothesis classes with uniformly bounded $\gamma$ (sequential) fat-shattering dimension for all $\gamma>0$, then ${\mathbb{E}}{\mathcal{F}}$ has finite $\gamma$ (sequential) fat-shattering dimension for all $\gamma>0$.
Proof  We prove this first for fat-shattering dimension, and then consider the sequential version.
If the class ${\mathbb{E}}{\mathcal{F}}$ has infinite $\gamma$ fat-shattering dimension for some $\gamma>0$, this is witnessed by a countable subset $Y^{\prime}\subseteq Y$. If we restrict the parameters of every class in ${\mathcal{F}}$ to $Y^{\prime}$, the resulting measurable family ${\mathcal{F}}^{\prime}$ is strongly measurable. Thus ${\mathbb{E}}{\mathcal{F}}^{\prime}$ having infinite $\gamma$ fat-shattering dimension implies, by Theorems 22 and and the connection between finite fat-shattering dimensions and Rademacher mean width, that there must be some $\delta>0$ for which the classes ${\mathcal{H}}\in{\mathcal{F}}^{\prime}$ have unbounded fat-shattering dimension, and the same applies to ${\mathcal{F}}$.
For the sequential version, the argument is the same, but citing Theorem 37.
However, Theorems 15 and 34 are still subject to the measurability assumptions of Facts 11 and 19 respectively.
Appendix D More detail on Fact 23
Recall that in the body we stated Fact 23, about passing from a uniformly stable family to stability of its expectation class. We cited Ben Yaacov and Keisler (2009); Ben Yaacov (2013) for this, and mention that the language used in these works is quite different. We now explain how to translate from the setting of these works to Fact 23.
Our statement is that if a measurable family ${\mathcal{F}}$ of hypothesis classes on $X$ parameterized by $Y$ is uniformly stable, then the expectation class ${\mathbb{E}}{\mathcal{F}}$ is also stable. The statements in Ben Yaacov and Keisler (2009); Ben Yaacov (2013) show that stability, as a property of a formula in a metric structure of continuous logic, is preserved under a construction known as “randomization”. For the definitions of continuous logic and metric structure, see Ben Yaacov (2013); Ben Yaacov and Keisler (2009) and references therein.
Given a measurable family ${\mathcal{F}}$ as above, we will construct such a metric structure, and check that stability of ${\mathbb{E}}{\mathcal{F}}$ follows from stability of a corresponding formula in the randomization structure.
In order to state this properly, we need to explain what it means for a formula to be stable in a metric structure or theory of continuous logic. We define this in terms of hypothesis classes to match the rest of this paper, but the resulting definition matches (Ben Yaacov and Usvyatsov, 2010, Definition 7.1).
Definition 70
In any metric structure $\mathcal{M}$, given a formula $\phi(x;y)$ of continuous logic, let ${\mathcal{H}}_{\mathcal{M},\phi}$ be the class on the $x$-sort of $\mathcal{M}$ parameterized by the $y$-sort of $\mathcal{M}$ and given by $({\mathcal{H}}_{\mathcal{M},\phi})_{y}(x)=\phi(x,y)$.
We say that $\phi(x;y)$ is stable in $\mathcal{M}$ when ${\mathcal{H}}_{\mathcal{M},\phi}$ is. We say that $\phi(x;y)$ is stable in a theory $T$ when it is stable in every model $\mathcal{M}\vDash T$.
We begin by recalling some definitions and results related to randomization from Ben Yaacov (2013). The special case of concept classes is worked out in Ben Yaacov and Keisler (2009).
Given a theory $T$ of continuous logic in the language $\mathcal{L}$, there is a language $\mathcal{L}_{R}$, known as the randomization of $\mathcal{L}$, and an $\mathcal{L}_{R}$-theory $T_{R}$ of continuous logic, known as the randomization of $T$, such that for every $\mathcal{L}$-formula $\phi(\bar{x})$ in the language of $T$, there is a corresponding $\mathcal{L}_{R}$-formula $\mathbb{E}[\phi(\bar{x})]$ in the language of $T_{R}$.
We defer to Ben Yaacov (2013) for the exact construction of $\mathcal{L}_{R}$ and $T_{R}$. Below we summarize their most relevant properties:
Definition 71 (See Ben Yaacov (2013, Definitions 3.4, 3.23))
Let $(\Omega,\Sigma,\mu)$ be a probability space, and let $(\mathcal{M}_{\omega}:\omega\in\Omega)$ be a family of metric structures in a relational language of continuous logic. We call $(\mathcal{M}_{\omega}:\omega\in\Omega)$ a random family of structures when for all relation symbols $R(\bar{x})$ in this language, and tuples $\bar{a}$ of random variables $a:\omega\to\bigcup_{\omega}\mathcal{M}_{\omega}$ where $a(\omega)\in\mathcal{M}_{\omega}$, the function $\omega\mapsto R(\bar{a}(\omega))$ is measurable.
Fact 27 (Ben Yaacov (2013, Corollary 3.24))
Let $T$ be a theory of continuous logic, and let $(\mathcal{M}_{\omega}:\omega\in\Omega)$ be a random family of models of $T$. Then there is a metric $\mathcal{L}_{R}$-structure $\mathcal{R}\vDash T_{R}$, consisting of the random variables $a:\omega\to\bigcup_{\omega}\mathcal{M}_{\omega}$ where $a(\omega)\in\mathcal{M}_{\omega}$, where for every formula $\phi(\bar{x})$,
$(\mathbb{E}[\phi(\bar{a})])^{\mathcal{R}}=\mathbb{E}_{\mu}\left[\phi(\bar{a}(\omega)))^{\mathcal{M}_{\omega}}\right].$
Fact 28 (Ben Yaacov (2013, Theorem 4.9))
If $\phi(\bar{x})$ is stable in $T$, then $\mathbb{E}[\phi(\bar{x})]$ is stable in $T_{R}$.
Recall that ${\mathcal{F}}=({\mathcal{H}}_{\omega}:\omega\in\Omega)$ consists of hypothesis classes ${\mathcal{H}}_{\omega}$ on $X$ parameterized by $Y$, where $(\Omega,\Sigma,\mu)$ is an atomless probability space with respect to which each function $\omega\mapsto{\mathcal{H}}_{\omega}(x,y)$ with $x\in X,y\in Y$ is measurable.
We now construct a measurable family of structures based on ${\mathcal{H}}_{\omega}$. Given $\omega\in\Omega$, let $\mathcal{M}_{\omega}$ be a metric structure with two sorts: $X,Y$. To each sort we assign the discrete metric ($d(x,y)=1$ when $x\neq y$), so all functions on these sorts will be uniformly continuous. Then we make this a structure in a language with one relation symbol, $\phi(x,y)$, where $x$ is a variable of sort $X$ and $y$ is a variable of sort $Y$. We interpret this symbol in $\mathcal{M}_{\omega}$ by $\phi(x,y)=({\mathcal{H}}_{\omega})_{y}(x)$.
Now let $T$ be the shared theory of all the structures $\mathcal{M}_{\omega}$.
Lemma 72
The partitioned formula $\phi(x;y)$ is stable in the theory $T$.
Proof  This means that for every $\mathcal{M}\vDash T$, the class ${\mathcal{H}}_{\mathcal{M},\phi}$ on the $X$-sort of $\mathcal{M}$ parameterized by the $Y$-sort of $\mathcal{M}$ and given by $({\mathcal{H}}_{\mathcal{M},\phi})_{y}(x)=\phi(x,y)$ is stable.
Note that the class ${\mathcal{H}}_{\mathcal{M}_{\omega},\phi}$ is just ${\mathcal{H}}_{\omega}$.
Because $T$ is the common theory of the structures $\mathcal{M}_{\omega}$, and for each $r<s$, the $(r,s)$-threshold dimensions of the classes ${\mathcal{H}}_{\mathcal{M}_{\omega},\phi}$ are uniformly bounded, the same bound is implied by some condition proven by the theory $T$, so this same bound also holds for any other $\mathcal{M}\vDash T$, so $\phi(x;y)$ is stable in $T$.
This implies also that $\mathbb{E}[\phi(x,y)]$ is stable in $T_{R}$. This allows us to construct stable hypothesis classes from models of $T_{R}$. We see that there is a model $\mathcal{R}\vDash T_{R}$ consisting of random variables where for each pair $a,b$ of random variables in the $X$-sort and $Y$-sort respectively,
$(\mathbb{E}[\phi(a,b)])^{\mathcal{R}}=\mathbb{E}_{\mu}\left[\phi(a(\omega),b(\omega)))^{\mathcal{M}_{\omega}}\right].$
We now claim that the hypothesis class ${\mathbb{E}}{\mathcal{F}}$ embeds into the class defined by $\mathbb{E}[\phi(x,y)]$ in this structure, and is thus stable. To see this, note that for every $x\in X$, $y\in Y$, the constant random variables $a,b$ taking values $x,y$ respectively are elements of the appropriate sorts of $\mathcal{R}$, by the construction in Ben Yaacov (2013). Thus for these elements of $\mathcal{R}$,
$(\mathbb{E}[\phi(a,b)])^{\mathcal{R}}=\mathbb{E}_{\mu}\left[\phi(x,y)^{\mathcal{M}_{\omega}}\right]=\mathbb{E}_{\mu}\left[({\mathcal{H}}_{\omega})_{y}(x)\right]={\mathbb{E}}{\mathcal{F}}_{y}(x).$